OpenStudy (anonymous):

An equation of a line through(6,11) which is perpendicular to the line y=3x+3 has a slope of? and a y intercept of?

OpenStudy (anonymous):

if you know a line is perpendicular to some other line with a known slope, you can easily get the slope of the line perpendicular... just take the opposite reciprocal of the known slope.

OpenStudy (anonymous):

what's the slope of the line given?

OpenStudy (anonymous):

what's the slope of the line y=3x+3

OpenStudy (anonymous):

(6, 11) is the coordinate of a point. this is a point. have you seen this equation before? : y = mx + b

OpenStudy (anonymous):

what does the m represent?

OpenStudy (anonymous):

that number, m, is your slope. so we know the slope of y = 3x + 3. the slope is 3. so according to what i stated earlier, what is the slope that is perpendicular to slope 3?

OpenStudy (anonymous):

take the opposite, reciprocal of 3....

OpenStudy (anonymous):

the perpendicular has slope -1/3... you now have the slope that's perpendicular to the given line and you have a point that goes through the point you want at (6, 11) apply this to y-y0 = m(x-x0) for the equation of the line you want... (x0, y0) is the point the line goes through.